OpenStudy (anonymous):

Simplify and state the restriction 12x^4-25x+12/3x^2+2x-8

OpenStudy (anonymous):

Is this your question? \(\Large{\frac{12x^4-25x+12}{3x^2+2x-8}}\)

OpenStudy (anonymous):

If so, both the numerator and denominator need to be factored first.

OpenStudy (anonymous):

How do you do that?

OpenStudy (anonymous):

I use \(\LaTeX\) because I'm already familiar with the language, but you can use the Equation editor button under the reply box. \(\Large{\frac{12x^4-25x+12}{3x^2+2x-8}}\) `\(\Large{\frac{12x^4-25x+12}{3x^2+2x-8}}\)` The \Large{ } just makes all the characters bigger and more readable

OpenStudy (anonymous):

So what do you do next then?

OpenStudy (anonymous):

You need to factor the numerator, then factor the denominator.

OpenStudy (anonymous):

\(12x^4-25x+12\) First you need two numbers that will multiply to be \(12x^4\)

OpenStudy (anonymous):

Sol like 4x^2 and 3x^2?

OpenStudy (anonymous):

That could be. Remember that 12 has factors 1,12 and 2,6 and 3,4 - It could be any one of those. Have you learned the quadratic formula?

OpenStudy (anonymous):

Not yet :/

OpenStudy (anonymous):

recheck the problem and make sure i have it written exactly correct. Check the coefficients, exponents, numbers and signs.

OpenStudy (anonymous):

I'm hoping the \(12x^4\) should be \(12x^2\).

OpenStudy (anonymous):

Ohh you are right it is 12x^2 im so sorry about that

OpenStudy (anonymous):

Great. Factor the denominator and we will come back to the numerator.

OpenStudy (anonymous):

So is it 3x-16 times 4x-9?

OpenStudy (anonymous):

FOIL your factors to check.

OpenStudy (anonymous):

\(3x^2+2x-8\) \((3x\pm?)(x\pm?)\) Remember that to get your middle term, you will combine the products of the inside and outside terms.

OpenStudy (anonymous):

Your last terms have to multiply to be -8 so your choices there are 1,8 and 2,4 and 4,2 and 8,1